OpenStudy (anonymous):

How do the graphs of y = 1/x and y = 5/x+6 compare?

OpenStudy (anonymous):

i just need help doing it, i don't just want the answer

OpenStudy (defiance):

Hello!

OpenStudy (anonymous):

hi! can you help?

OpenStudy (defiance):

Sure thing.

OpenStudy (defiance):

Do you have any ideas? Or should we start from scatch?

OpenStudy (anonymous):

scratch?

OpenStudy (defiance):

Okay.

OpenStudy (defiance):

First look at y = 1/x if x is 0 then y= 1/0 Which is undefined. But think like this: instead of 0, let x be a small number like 1/100000 when we divide 1/ (1/100000) we get 100000 (a big number) the point is as you get close to x=0 y gets really big when x=1, y =1/1 = 1 when x is really big y = 1/10000 which is close to 0 so here are 3 points (0, infinity), (1,1), and (infinity, 0)

OpenStudy (defiance):

Look at it like this now, \[ y = \frac{5}{x+6} \]

OpenStudy (defiance):

Got it so far?

OpenStudy (defiance):

\[ y = \frac{5}{x+6} \] at a certain x value x+6 = -6

OpenStudy (anonymous):

i saw this explanation on another post but its kind of confusing. do i have to use 100000

OpenStudy (defiance):

When is y= 1 (to pick a point) 1= 5/(x+6) Multiply both sides by x+6 to get x+6= 5 x= -1 so (-1,1) is on the graph then, when x gets really big y = 5/big number ---> 0

OpenStudy (defiance):

If you still need more help: http://www.regentsprep.org/Regents/math/algtrig/ATP9/funclesson1.htm

OpenStudy (anonymous):

Is the second equation \[y = \dfrac{5}{x}+6 \] or \[y = \dfrac{5}{x+6} \]